# Photodriven Mott insulating heterostructures: A steady-state study of impact ionization processes

### v1

@@ I Introduction
In the present manuscript we characterize a prototype of solar cell based on the heterostructure schematically depicted in Fig. 1 in terms of the occurrence of II. The system under investigation consists of $\text{L}=4$ Mott insulating layers located between two wide-band metallic leads which act as charge collectors to harvest the energy. The Mott region is separated from the metallic leads by two narrow-band layers with energies aligned with the upper and lower Hubbard bands as in Fig. 2, which represent the oxide contact layers. An external time-periodic electric field induces a Floquet steady state with a current flowing from the left to the right side of the heterostructure, against the natural gradient established by the leads’ chemical potentials. Due to the polar interfaces located between the insulating layers [12], an internal electric field builds up in the heterostructure, acting as a separation mechanism for the electron-hole (e-h) pairs produced by the periodic driving.
@@
Our results indicate that II, which we link to the presence of a clear increase (“kink”) in the photocurrent for driving frequency strengths around twice the gap, plays an important role when the correlated region is connected to narrow-band contacts. This sudden increase in the photocurrent is generally supported by a corresponding kink in the double occupation and its fluctuation around its mean value (see Eq. (20) for details) as well as by a non-negligible occupation of the upper bands for driving frequencies compatible with II. Upon decreasing the electric field amplitude $E_{0}$, our results suggest that II is accompanied by a change of behavior of the photocurrent from quadratic to linear as function of $E_{0}$. Factors that hinder II turn out to be (i) a stronger hybridization to the leads, as we already observed in the case of a single layer [32], as well as (ii) a direct connection to wide-band metallic leads without intermediate narrow-band contacts.
@@ III.3 Physical quantities
In order to characterize the spectral properties of the system at hand we study the local electron density of states (DOS) $A_{z}(\omega)$ and occupation function $N_{z}(\omega)$ which are defined as
@@ IV Results
Figure 2: Schematic representation of the relevant processes occurring in the system considered in this manuscript. Electrons are injected into the correlated region (green) from the (almost full) left reservoir and drained into the (almost empty) right one (red). The potential drop across the correlated region is denoted as $\Phi$ and is a linear function of the coordinate $z$ of the layers. We set $\mu_{r}>\mu_{l}$ so that photovoltaic energy is collected from the device when current flows from left to right. (a) Direct excitation processes. When $\Omega<2\Delta_{\text{pg}}$, the photoexcited electron in the upper band escapes directly into the right lead by tunneling through the upper bands. (b) Impact ionization processes. For large values of $\Omega$, i.e. $\Omega\geq 2\Delta_{\text{pg}}$, an electron promoted to the upper band can excite, in turn, a second electron across the pseudogap by transferring energy to it via electron-electron scattering, so that two carriers per photon can now tunnel towards the right lead.
The chemical potentials are chosen such that $\mu_{\text{r}}=-\mu_{\text{l}}>0$, so the energy is harvested from periodic driving when an electron current flows from the left to the right lead, against the tendency established by the gradient given by the leads’ chemical potentials. Due to the hybridization with the leads, the local DOS of the correlated layers does not exhibit a clear gap, but rather only a pseudogap $\Delta_{\text{pg}}\approx 3$, see also Ref. [32], while the leads’ bandwidth $W_{\text{b}}\approx 8.5$.
In order for II processes to be energetically allowed, the bandwidth of the upper Hubbard band (UHB), which in our case is roughly equal to the leads’ DOS, has to be at least twice the size of the pseudogap $\Delta_{\text{pg}}$, i.e. $W_{\text{b}}\geq 2\Delta_{\text{pg}}$. Only in this case an electron promoted to the UHB has sufficient energy to excite a second one across the gap [19, 32].
@@ IV.1 Impact ionization
Figure 3: (a) Photocurrent $j$ [Eq. (16)] as function of the driving frequency $\Omega$ with error bars corresponding to $\sigma_{j}$. (b) Double occupation $N_{\text{D},z}$ as a function of the driving frequency $\Omega$ for all the layers of the correlated region. (c) Corresponding fluctuation $\Delta N_{\text{D},z}$ as function of the driving frequency $\Omega$ for the layers $z=1$ and $z=2$. The black arrows in (a)-(c) highlight the change in slope $\sim 2\Delta_{\textrm{pg}}$. Due to the PhI symmetry (see App. B), $\Delta N_{\text{D},\text{L}+1-z}=\Delta N_{\text{D},z}$ so the curves for $z=3$ and $z=4$ in (c) are omitted. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
The main purpose of this section is to study the onset of the II processes: for this reason we evaluate the photocurrent as function of the driving frequency $\Omega$. As argued in Ref. [19, 32], the onset of II can be identified by an appreciable change in slope (kink) in the $j$-$\Omega$ curve at $\Omega\sim 2\Delta_{\textrm{pg}}$, as it hints at an increased number of carriers being promoted to the UHB.
@@
  • For $\Omega<\Delta_{\text{pg}}$, we expect a negligible photocurrent produced by the residual DOS in the pseudogap.
  • For $\Delta_{\text{pg}}<\Omega<\Delta_{\text{pg}}+2W_{\text{b}}$ as shown in Fig. 2(a), an electron from a lower Hubbard band (LHB) gets photoexcited to the corresponding UHB and moves to the right along the potential drop without further exciting e-h pairs. Such process is often referred to as direct excitation (DE).
  • For $2\Delta_{\text{pg}}<\Omega<\Delta_{\text{pg}}+2W_{\text{b}}$ as shown in Fig. 2(b), a photoexcited electron in the UHB may excite a second one across the pseudogap via II, before both move to the right due to the potential drop.
  • For $\Omega>\Delta_{\text{pg}}+2W_{\text{b}}$, a photoexcited electron ends up near the border of the UHB where the DOS is greatly reduced, and the photocurrent gets strongly suppressed.
@@ IV.1.1 Photocurrent and spectral features
By the analysis of the $j$-$\Omega$ curve in Fig. 3(a) we see an increase of the photocurrent as $\Omega$ grows larger, with the maximum reached at $\Omega\approx 11$, which is followed by a decrease until $\Omega\approx 17$  [^6] The current $j$ reaches its minimum at $\Omega\approx 17$ because the effective current energy window given by the overlap of the layers’ UHBs is smaller than $\Delta_{\text{pg}}+2W_{\text{b}}$. Moreover, $j$ does not go to zero as one would expect because of the background current as shown in Ref. [19, 32].. A slight change of slope of the curve around $\Omega\approx 7$ seems to indicate an onset of II processes, which are indeed expected to start showing up at $\sim 2\Delta_{\text{pg}}$. This is corroborated by a similar behavior of the double occupation $N_{\text{D},z}$ as function of $\Omega$ for the different layers $z=\left\{1,\dots,4\right\}$ displayed in Fig. 3(b). All these curves show a change in slope around $\Omega\approx 7$, although this appears relatively weak for the $z=3$ and $4$ layers. This behavior is due to the different occupation in the layers as discussed in Sec. III.3. In order to concentrate on electronic correlations effects we plot the corresponding fluctuation $\Delta N_{\text{D},z}$ (see Eq. (20)) in Fig. 3(c) for $z=1$, $2$. These curves are much less $z$-dependent and, more importantly, display a clear kink at $\Omega\approx 7$, corroborating the onset of II beyond that driving frequency.
@@
Such indications of II processes are supported by Fig. 4, in which the electron DOS $A_{z}(\omega)$ and spectral occupation $N_{z}(\omega)$ are shown for selected values of the driving $\Omega$. As schematically represented in Fig. 2, we first notice that the electron DOS of all layers features a pseudogap for both $\Omega=5$ and $\Omega=11$.
Figure 4: Electron DOS $A_{z}(\omega)$ (solid lines) and occupation function $N_{z}(\omega)$ (shaded areas) for the different layers at (a) $\Omega=5$ and (b) $\Omega=11$. Dashed vertical lines in (a) mark the position of the leads’ chemical potentials $\mu_{\text{l}}$ and $\mu_{\text{r}}$, while those in (b) highlight the separation between the center of the bands of the layers $z=1$ and $z=4$, which roughly equals the potential drop $\Phi$, depicted in Fig. 2. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
At $\Omega=11$, where II is expected to happen, we observe that an important fraction of the electrons now occupies the UHB of all layers in the correlated region, see Fig. 4(b). As argued at the beginning of Sec. IV.1, a larger occupation of the UHB is consistent with an increased scattering probability among electrons across the pseudogap, the key mechanism for II.
@@ IV.2 Dependence on the electric field amplitude
In this section we investigate the behavior of the photocurrent and of II upon reducing the amplitude $E_{0}$ of the driving electric field  [^7] We point out that the field strength we are considering are rather in the range achieved with intense ultrashort pulsed lasers (see also [35]) and far from the intensity of solar radiation..
@@
Figure 5: Photocurrent $j$ (with error bars $\sigma_{j}$) as function of $E_{0}$ at (a)-(d) $\Omega=5$, (b)-(e) $\Omega=8$ and (c)-(f) $\Omega=11$ for different number of correlated layers L and hybridizations $v$. Horizontal dashed lines denote the value of the background current obtained with $E_{0}=0$ and refer to the curves of the same color. Green curves in (e)-(f) corresponding to $\text{L}=4$ and $v=0.4$ from (b)-(c) are replotted for comparison. Default parameters for the $\text{L}=4$ curves are specified in Table 1, while the $\text{L}=1$ curve is obtained with the same parameters used in Ref. [32]. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$, $W_{\text{b}}\approx 8.5$)
@@
These results suggest a correlation between the occurrence of II and the behavior of the $j$-$E_{0}$ curves. A rough explanation for the deviation from a quadratic behavior in the II case may be provided by the following argument. One expects the DE photocurrent to be proportional to $|E_{0}|^{2}$ for small $E_{0}$, while in presence of II $j\propto|E_{0}|^{2}[1+r_{\textrm{II}}]$, in which the II scattering rate $r_{\textrm{II}}$  [^9] The II scattering rate $r_{\textrm{II}}$ is proportional to $U^{2}$ for small $U$, with a probable saturation for large $U$. accounts for the creation of two doublons and two holons (see Ref. [1] for details), increases as function of $\Omega$ and contributes only when $\Omega\gtrsim 2\Delta_{\textrm{pg}}$. $r_{\textrm{II}}$ is certainly dependent on the occupation of the UHB, and may be responsible for the linear behavior of the photocurrent above a certain threshold field $E\geq E_{0,\textrm{th}}$. However, this behavior can only be deduced a posteriori from our results in Fig. 5. We argue that this crossover from a quadratic to a linear-like $j$-$E_{0}$ behavior may be considered as a signature of II. This argument is corroborated by the fact that for $E_{0}$-values slightly bigger than the supposed threshold field $E_{0,\textrm{th}}\sim 0.5$ the occupation functions of the UHB for $\Omega=8$ and especially $\Omega=11$ are clearly larger than for $\Omega=5$, as can be seen in Fig. 6  [^10] We display in Fig. 6 the spectral features only for $z=2$ because they are really similar for the other layers of the heterostructure..
Figure 6: Electron DOS $A_{z}(\omega)$ (dashed lines) and occupation function $N_{z}(\omega)$ (solid lines) relative to layer $z=2$ at (a) $E_{0}=0.75$ and (b) $E_{0}=1$ for different values of the driving frequency $\Omega$. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
The value of $E_{0,\textrm{th}}$, above which the curve $j$-$E_{0}$ becomes linear, increases when decreasing the number of layers $L$ in the central region, as one sees from Fig. 5(a)-(c), which displays the $j$-$E_{0}$ curves for $\text{L}=1$  [^11] The data for the $\text{L}=1$ case are obtained with the same parameters used in Ref. [32].. This can be understood by the observation that excited doublons exhibit a swifter departure from a more confined correlated region, which in turn decreases the probability of II for a fixed $E_{0}$. A similar effect is observed when increasing the hybridization to the leads as displayed in Fig. 5(e)-(f), as already observed in Ref. [32]. Due to the larger value $v=0.8$  [^12] The data for the $v=0.8$ are obtained with the default parameters specified in Table 1. which disfavors II, the linear regime sets in at larger field amplitudes ($E_{0}\gtrsim 1$) and it is more marked at larger driving frequencies ($\Omega=11$) with respect to $v=0.4$, see Fig. 5(e)-(f)). Support of this general trend, whereby II is disfavored by a larger hybridization and a smaller number of layers $L$ is given by Fig. 7  [^13] We choose in Fig. 7 to plot th the spectral features for $z=1$ for the $L=4$ setups for consistency with the $L=1$ case. The other layers of the correlated region show really close characteristics.. Here the occupation of the UHB is significantly reduced for $L=1$ and $v=0.8$ with respect to $v=0.4$, indicating a suppression of II.
Figure 7: Electron DOS $A_{z}(\omega)$ (dashed lines) and occupation function $N_{z}(\omega)$ (solid lines) relative to layer $z=1$ at (a) $E_{0}=0.75$ and (b) $E_{0}=1$ for different number of correlated layers L and hybridizations $v$ at $\Omega=11$. Default parameters for the $\text{L}=4$ curves are specified in Table 1, while the $\text{L}=1$ curve is obtained with the same parameters used in Ref. [32]. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$, $W_{\text{b}}\approx 8.5$)
@@
Figure 8: (a) Photocurrent $j$ and (b) fluctuation of double occupancy $\Delta N_{\text{D},z}$ for the layer $z=2$, as function of the driving frequency $\Omega$ for selected values of the electric field amplitude $E_{0}$. Error bars corresponding to $\sigma_{j}$ in (a) are not shown for better visualization of the data. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
To complete our analysis, in Fig. 8 we analyze the behavior of the $j$-$\Omega$ curves for values of the field amplitude $E_{0}=\left\{0.75,\dots,1.75\right\}$. These curves display no appreciable change in slope for $E_{0}\lesssim 1.25$ and only a slight bend for larger fields (see Fig. 8(a)) around $\Omega\approx 5-7$, even though II is expected to occur at $E_{0}\gtrsim 0.5$, as suggested by Fig. 5(b)-(c)). On the other hand, the $\Delta N_{\text{D}}$-$\Omega$ curve in Fig. 8(b) displays a small but clearer kink already at smaller field amplitudes, in support of II. The behavior of this quantity as function of $E_{0}$ in Fig. 9 confirms the II onset from $E_{0}\gtrsim 0.5$, with a clear transition from the small $E_{0}$ quadratic regime to a linear one for $\Omega=8,11$ (green curve) in accordance with Fig. 5. Also the curves for $L=1$ and $v=0.8$ in Fig. 9(b)-(c) undergo a crossover to a linear behavior at larger $E_{0}$-values ($E_{0}\gtrsim 1$), here especially evident for $\Omega=11$, as already observed in Fig. 5.
Figure 9: Fluctuation of double occupancy $\Delta N_{\text{D},z}$ for the layer $z=1$ as function of $E_{0}$ at (a) $\Omega=5$, (b) $\Omega=8$ and (c) $\Omega=11$ for different number of correlated layers L and hybridizations $v$. Default parameters for the $\text{L}=4$ curves are specified in Table 1, while the $\text{L}=1$ curve is obtained with the same parameters used in Ref. [32]. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{pg}}\approx 3$, $W_{\text{b}}\approx 8.5$)

### v2

@@ I Introduction
In the present manuscript we characterize a prototype of solar cell based on the heterostructure schematically depicted in Fig. 1 in terms of the occurrence of II. The system under investigation consists of $\text{L}=4$ Mott insulating layers located between two wide-band metallic leads which act as charge collectors to harvest the energy. The Mott region is separated from the metallic leads by two narrow-band layers with energies aligned with the upper and lower Hubbard bands as in Fig. 2, which represent the oxide contact layers. An external time-periodic electric field induces a Floquet steady state with an average current flowing from the left to the right lead, against the potential energy established by the leads’ chemical potentials. Due to the polar interfaces located between the insulating layers [12], an internal electric field builds up in the heterostructure, acting as a separation mechanism for the electron-hole (e-h) pairs produced by the periodic driving.
@@
Our results indicate that II, which we link to the presence of a sharp increase (“kink”) in the photocurrent for driving frequency strengths around twice the gap, plays an important role when the correlated region is connected to narrow-band contacts. This sudden increase in the photocurrent is generally supported by a corresponding kink in the double occupation and its fluctuation around its mean value (see Eq. (20) for details) as well as by a non-negligible occupation of the upper bands for driving frequencies compatible with II. Upon decreasing the electric field amplitude $E_{0}$, our results suggest that II is accompanied by a change of behavior of the photocurrent from quadratic to linear as function of $E_{0}$. Factors that hinder II turn out to be (i) a stronger hybridization to the leads, as we already observed in the case of a single layer [32], as well as (ii) a direct connection to wide-band metallic leads without intermediate narrow-band contacts.
@@ III.3 Physical quantities
In order to characterize the spectral properties of the system at hand we study the local electron density of states (DOS) $A_{z}(\omega)$ and occupation function $N_{z}(\omega)$  [^4] The occupation function $N_{z}(\omega)$ indicates how many electronic states per spin per layer are occupied with energy $\omega$ in the layer $z$. The integral over the whole energy range of such quantity gives the number of particles per spin per layer $n_{z}$, i.e. $n_{z}=\int_{-\infty}^{+\infty}\mathop{}\!\mathrm{d}\omega N_{z}(\omega)$. which are defined as
@@ IV Results
Figure 2: Schematic representation of the relevant processes occurring in the system considered in this manuscript. Electrons are injected into the correlated region (green) from the (almost full) left reservoir and drained into the (almost empty) right one (red). The potential drop across the correlated region is denoted as $\Phi$ and is a linear function of the coordinate $z$ of the layers. We set $\mu_{r}>\mu_{l}$ so that photovoltaic energy is collected from the device when current flows from left to right. (a) Direct excitation processes. When $\Omega<2\Delta_{\text{g}}$, the photoexcited electron in the upper band escapes directly into the right lead by tunneling through the upper bands. (b) Impact ionization processes. For large values of $\Omega$, i.e. $\Omega\geq 2\Delta_{\text{g}}$, an electron promoted to the upper band can excite, in turn, a second electron across the gap by transferring energy to it via electron-electron scattering, so that two carriers per photon can now tunnel towards the right lead.
The chemical potentials are chosen such that $\mu_{\text{r}}=-\mu_{\text{l}}>0$, so the energy is harvested from periodic driving when an electron current flows from the left to the right lead, against the chemical potential difference. The gap of the correlated layers is $\Delta_{\text{g}}\approx 3$  [^7] Due to the hybridization with the leads, a finite spectral weight is present in the gap region of the correlated layers’ local DOS, see also Ref. [32]., while the leads’ bandwidth is $W_{\text{b}}\approx 8.5$.
In this manuscript our focus is on II processes, in which an electron, promoted to the upper Hubbard band by absorbing the energy of a photon $\Omega$, excites a second electron across the gap by exchanging energy to it via Coulomb interaction.In order for such processes to be energetically allowed, the bandwidth of the upper Hubbard band (UHB), which in our case is roughly equal to the leads’ DOS, has to be at least twice the size of the gap $\Delta_{\text{g}}$, i.e. $W_{\text{b}}\geq 2\Delta_{\text{g}}$. Only in this case an electron promoted to the UHB has sufficient energy to excite a second one across the gap [19, 32].
@@ IV.1 Impact ionization
Figure 3: (a) Photocurrent $j$ [Eq. (16)] as function of the driving frequency $\Omega$ with error bars corresponding to $\sigma_{j}$. (b) Numerical derivative of the photocurrent $\text{d}j/\text{d}\Omega$ as function of the driving frequency $\Omega$. (c) Fluctuation of the double occupation $\Delta N_{\text{D},z}$ as function of the driving frequency $\Omega$ for the layers $z=1$ and $z=2$. (d) Double occupation $N_{\text{D},z}$ as a function of the driving frequency $\Omega$ for all the layers of the correlated region. The black arrows in (a)-(b)-(c) highlight the change in slope $\sim 2\Delta_{\textrm{g}}$. Due to the PhI symmetry (see App. B), $\Delta N_{\text{D},\text{L}+1-z}=\Delta N_{\text{D},z}$ so the curves for $z=3$ and $z=4$ in (c) are omitted. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
The main purpose of this section is to study the onset of the II processes: for this reason we evaluate the photocurrent as function of the driving frequency $\Omega$. As argued in Ref. [19, 32], the onset of II can be possibly identified by an appreciable change in slope (kink) in the $j$-$\Omega$ curve at $\Omega\sim 2\Delta_{\textrm{g}}$, as it hints at an increased number of carriers being promoted to the UHB, which contribute to a steeper increase of the photocurrent.
@@
  • For $\Omega<\Delta_{\text{g}}$, we expect a negligible photocurrent produced by the residual DOS in the gap, due to the hybridization with the leads.
  • For $\Delta_{\text{g}}<\Omega<\Delta_{\text{g}}+2W_{\text{b}}$ as shown in Fig. 2(a), an electron from a lower Hubbard band (LHB) gets photoexcited to the corresponding UHB and moves to the right along the potential drop without further exciting e-h pairs. Such process is often referred to as direct excitation (DE).
  • For $2\Delta_{\text{g}}<\Omega<\Delta_{\text{g}}+2W_{\text{b}}$ as shown in Fig. 2(b), a photoexcited electron in the UHB may excite a second one across the gap via II, before both move to the right due to the potential drop.
  • For $\Omega>\Delta_{\text{g}}+2W_{\text{b}}$, a photoexcited electron ends up near the border of the UHB where the DOS is greatly reduced, and the photocurrent gets strongly suppressed.
@@ IV.1.1 Photocurrent and spectral features
By the analysis of the $j$-$\Omega$ curve in Fig. 3(a) we see an increase of the photocurrent as $\Omega$ grows larger, with the maximum reached at $\Omega\approx 11$, which is followed by a decrease until $\Omega\approx 17$  [^8] The current $j$ reaches its minimum at $\Omega\approx 17$ because the effective current energy window given by the overlap of the layers’ UHBs is smaller than $\Delta_{\text{g}}+2W_{\text{b}}$. Moreover, $j$ does not go to zero as one would expect because of the background current as shown in Ref. [19, 32].. A slight change of slope of the curve around $\Omega\approx 7$ seems to indicate an onset of II processes, which are indeed expected to start showing up at $\sim 2\Delta_{\text{g}}$, confirmed by the numerical derivative of $j$ with respect to $\Omega$ in Fig. 3(b). This is also corroborated by a similar behavior of the double occupation $N_{\text{D},z}$ as function of $\Omega$ for the different layers $z=\left\{1,\dots,4\right\}$ displayed in Fig. 3(d). All these curves show a change in slope around $\Omega\approx 7$, although this appears relatively weak for the $z=3$ and $4$ layers. This behavior is due to the different occupation in the layers as discussed in Sec. III.3. In order to concentrate on electronic correlations effects we plot the corresponding fluctuation $\Delta N_{\text{D},z}$ (see Eq. (20)) in Fig. 3(c) for $z=1$, $2$. These curves are much less $z$-dependent and, more importantly, display a clear kink at $\Omega\approx 7$, corroborating the onset of II beyond that driving frequency.
@@
Such indications of II processes are supported by Fig. 4, in which the electron DOS $A_{z}(\omega)$ and spectral occupation $N_{z}(\omega)$ are shown for selected values of the driving $\Omega$. As schematically represented in Fig. 2, we first notice that the electron DOS of all layers features a finite spectral weight in the gap for both $\Omega=5$ and $\Omega=11$.
Figure 4: Electron DOS $A_{z}(\omega)$ (solid lines) and occupation function $N_{z}(\omega)$ (shaded areas) for the different layers at (a) $\Omega=5$ and (b) $\Omega=11$. Dashed vertical lines in (a) mark the position of the leads’ chemical potentials $\mu_{\text{l}}$ and $\mu_{\text{r}}$, while those in (b) highlight the separation between the center of the bands of the layers $z=1$ and $z=4$, which roughly equals the potential drop $\Phi$, depicted in Fig. 2. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
At $\Omega=11$, where II is expected to happen, we observe that an important fraction of the electrons now occupies the UHB of all layers in the correlated region, see Fig. 4(b). As argued at the beginning of Sec. IV.1, a larger occupation of the UHB is consistent with an increased scattering probability among electrons across the gap, the key mechanism for II.
@@ IV.2 Dependence on the electric field amplitude
In this section we investigate the behavior of the photocurrent and of II upon reducing the amplitude $E_{0}$ of the driving electric field  [^9] We point out that the field strength we are considering are rather in the range achieved with intense ultrashort pulsed lasers (see also [35]) and far from the intensity of solar radiation. See also the discussion in Sec. V.
@@
Figure 5: Photocurrent $j$ (with error bars $\sigma_{j}$) as function of $E_{0}$ at (a)-(d) $\Omega=5$, (b)-(e) $\Omega=8$ and (c)-(f) $\Omega=11$ for different number of correlated layers L and hybridizations $v$. Horizontal dashed lines denote the value of the background current obtained with $E_{0}=0$ and refer to the curves of the same color. Green curves in (e)-(f) corresponding to $\text{L}=4$ and $v=0.4$ from (b)-(c) are replotted for comparison. Default parameters for the $\text{L}=4$ curves are specified in Table 1, while the $\text{L}=1$ curve is obtained with the same parameters used in Ref. [32]. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{g}}\approx 3$, $W_{\text{b}}\approx 8.5$)
@@
These results suggest a correlation between the occurrence of II and the behavior of the $j$-$E_{0}$ curves. A rough explanation for the deviation from a quadratic behavior in the II case may be provided by the following argument, deduced a posteriori from our results in Fig. 5. One expects the DE photocurrent to be proportional to $|E_{0}|^{2}$ for small $E_{0}$, while in presence of II $j\propto|E_{0}|^{2}[1+r_{\textrm{II}}]$, in which the II scattering rate $r_{\textrm{II}}$  [^11] The II scattering rate $r_{\textrm{II}}$ is proportional to $U^{2}$ for small $U$, with a probable saturation for large $U$. accounts for the creation of two doublons and two holons (see Ref. [1] for details), increases as function of $\Omega$ and contributes only when $\Omega\gtrsim 2\Delta_{\textrm{g}}$. $r_{\textrm{II}}$ is certainly dependent on the occupation of the UHB, and may be responsible for the linear behavior of the photocurrent above a certain threshold field $E\geq E_{0,\textrm{th}}$. We argue that this crossover from a quadratic to a linear-like $j$-$E_{0}$ behavior may be considered as a signature of II.
@@
The value of $E_{0,\textrm{th}}$, above which the curve $j$-$E_{0}$ becomes linear, increases when decreasing the number of layers $L$ in the central region, as one sees from Fig. 5(a)-(c), which displays the $j$-$E_{0}$ curves for $\text{L}=1$  [^12] The data for the $\text{L}=1$ case are obtained with the same parameters used in Ref. [32].. This can be understood by the observation that excited doublons exhibit a swifter departure from a more confined correlated region, which in turn decreases the probability of II for a fixed $E_{0}$. A similar effect is observed when increasing the hybridization to the leads as displayed in Fig. 5(e)-(f), as already observed in Ref. [32]. Due to the larger value $v=0.8$  [^13] The data for the $v=0.8$ are obtained with the default parameters specified in Table 1. which disfavors II, the linear regime sets in at larger field amplitudes ($E_{0}\gtrsim 1$) and it is more marked at larger driving frequencies ($\Omega=11$) with respect to $v=0.4$, see Fig. 5(e)-(f)).
@@
Figure 6: (a) Photocurrent $j$, (b) numerical derivative of the photocurrent $\text{d}j/\text{d}\Omega$ and (c) fluctuation of double occupancy $\Delta N_{\text{D},z}$ for the layer $z=2$, as function of the driving frequency $\Omega$ for selected values of the electric field amplitude $E_{0}$. Error bars corresponding to $\sigma_{j}$ in (a) are not shown for better visualization of the data. Default parameters are specified in Table 1. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
To complete our analysis, in Fig. 6 we analyze the behavior of the $j$-$\Omega$ curves for values of the field amplitude $E_{0}=\left\{0.75,\dots,1.75\right\}$. These curves display no appreciable change in slope for $E_{0}\lesssim 1.25$ and only a slight bend for larger fields (see Fig. 6(a)) around $\Omega\approx 5-7$, even though II is expected to occur at $E_{0}\gtrsim 0.5$, as suggested by Fig. 5(b)-(c)). On the other hand, the numerical derivative of the photocurrent $\text{d}j/\text{d}\Omega$ in Fig. 6(b) increases for $\Omega\sim 2\Delta_{\textrm{g}}$ and the $\Delta N_{\text{D}}$-$\Omega$ curve in Fig. 6(c) displays a small but clearer kink already at smaller field amplitudes, in support of II. The behavior of this quantity as function of $E_{0}$ in Fig. 7 confirms the II onset from $E_{0}\gtrsim 0.5$, with a clear transition from the small $E_{0}$ quadratic regime to a linear one for $\Omega=8,11$ (green curve) in accordance with Fig. 5. Also the curves for $L=1$ and $v=0.8$ in Fig. 7(b)-(c) undergo a crossover to a linear behavior at larger $E_{0}$-values ($E_{0}\gtrsim 1$), here especially evident for $\Omega=11$, as already observed in Fig. 5.
Figure 7: Fluctuation of double occupancy $\Delta N_{\text{D},z}$ for the layer $z=1$ as function of $E_{0}$ at (a) $\Omega=5$, (b) $\Omega=8$ and (c) $\Omega=11$ for different number of correlated layers L and hybridizations $v$. Default parameters for the $\text{L}=4$ curves are specified in Table 1, while the $\text{L}=1$ curve is obtained with the same parameters used in Ref. [32]. (Here $\Phi=2$, $\mu_{\text{l}}=-1$, $\Delta_{\text{g}}\approx 3$, $W_{\text{b}}\approx 8.5$)